# pH-Driven Fluorescence in Solid-State Lysine

### v1

@@ Abstract
Critical aggregation concentration (CAC) value for Lys in water was determined by fluorescence spectroscopy. In detail, two different fluorophores: a) 8-anilino-1-naphthalene sulfonic acid ammonium salt (ANS) and b) pyrene (pyr) were separately titrated with increasing amounts of the amino acid water solution. Fluorescence spectra were collected at room temperature on a Jasco spectrofluorophotometer (Model FP-8550), acquiring measurements in a quartz cell with a 1.0 cm path length at room temperature. The other experimental settings are: excitation and emission bandwidths = 5 nm; recording speed = 120 nmmin-1, excitation wavelength = 350 nm and 335 nm for ANS and pyr, respectively. The measurement was performed by adding small aliquots of the amino acid in 200 L of 200 M ANS or 10 M pyr water solution. At the end of the titration, the blank was subtracted. Fluorescence spectra were corrected for the blank and adjusted for the dilution. Tendency lines were extracted using a least squares method extrapolation. Titrations were conducted in triplicate. In the ANS method, the CAC value was determined by linear least-squares fitting the fluorescence emission at 475 nm, upon excitation at 350 nm versus the amino acid concentration as previously reported58. In the pyr method, the CAC value was determined by the break-point of the fitting of I1/I3 upon excitation at 335 nm versus the amino acid concentration59.
S2 Computational Details
S2.1 Classical molecular dynamics
Lys systems corresponding to different protonation states were modeled using crystallographic structures of L-Lysine sulfate (LLS), L-Lysine monohydrochloride (LLMHCl), and L-Lysine hemihydrate (LLH). Supercells of appropriate size were built for each system: 1588 for LLS, 11511 for LLMHCl, and 7116 for LLH. Classical molecular dynamics (MD) simulations were performed using the AMBER MD package version 202360, 61 with the Generalized Amber Force Field (GAFF) 62 applied to all molecular mechanics (MM) atoms. Each system was first equilibrated in the NVT ensemble at 300 K for 10 ns using a 2 fs time step and a Langevin thermostat63 with a collision frequency, , of ps-1. Next, a 50 ns production run was performed under the same conditions, with all bonds involving hydrogen atoms constrained. From this trajectory, 50 independent frames were extracted at 1 ns intervals and used as starting geometries for the subsequent quantum mechanics/molecular mechanics (QM/MM) simulations.
S2.2 Ground state QM/MM
To accurately capture the luminescent properties of the Lys crystals, it is essential to account for the influence of the surrounding environment. As fully QM treatment of the entire system is computationally prohibitive, we employed a hybrid QM/MM approach64, 65. For each crystal, we selected relevant QM regions (dimers; see Fig. 4a–c in the Main Text) from the 50 uncorrelated frames extracted from classical MD simulations. Thermal equilibration of each QM region at 300 K was performed using ORCA 5.0.466, 67, 68, 69, 70 (QM) interfaced with AMBER 202360, 61 (MM). The QM region was treated using the CAM-B3LYP exchange-correlation (xc) functional71 and the def2-SVP basis set72, 73 with the RIJ-COSX approximation74 to accelerate Coulomb and exact exchange integral evaluations. Equilibration was carried out for 1 ps in the NVT ensemble at 300 K using a Langevin thermostat63 with tight coupling ( ps-1) and no bond constraints. A 0.5 fs time step was used for 1 ps, followed by an additional 1 ps with a standard thermostat coupling ( ps-1). Final nuclear coordinates and velocities from these equilibrated QM/MM trajectories were used as initial conditions for the non-adiabatic molecular dynamics (NAMD) simulations described below.
In addition, vertical excitation energies were computed from these equilibrated structures to characterize the absorption spectra and identify relevant transitions. These time-dependent DFT (TDDFT) calculations were performed using the same ORCA/AMBER QM/MM framework, but with no MD propagation (0 MD steps), and the same xc functional, basis set, and approximations described above. For each Lys crystal system, the ten lowest singlet excited states were computed for 50 equilibrated initial conditions. Excitation plots shown in Fig. 4e in the Main Text correspond to the averaged energies of the lowest-energy absorption transition across these 50 frames. Molecular orbitals involved in these transitions were obtained using a post-processing tool available in ORCA.
S2.3 Non-adiabatic molecular dynamics
NAMD simulations were performed using the Trajectory Surface Hopping (TSH) method, specifically employing the SHARC formalism75, 76, which extends Tully’s Fewest Switches Surface Hopping (FSSH)77 to incorporate both time-dependent non-adiabatic couplings (TD-NACs) and spin–orbit couplings on an equal footing. In this approach, nuclei evolve classically on quantum potential energy surfaces computed on-the-fly. For further background, we refer the reader to comprehensive reviews of TSH methodology.78, 79, 80, 81.
All simulations–50 for each system–included the first 5 singlet states () to adequately describe the lowest excited states relevant for fluorescence studies. Each trajectory was initialized on the first excited state () and propagated in the NVE ensemble for 500 fs with a time-step of 0.5 fs using the velocity Verlet integration algorithm82, 83. The functional and basis set used in the ground-state QM/MM dynamics were kept unchanged in the NAMD simulations. Excitation energies and oscillator strengths were calculated at the TDDFT84 level, within the Tamm-Dancoff approximation (TDA)85, which improves TDDFT behavior near conical intersections86, 87, 88. NACs were computed via the Hammes-Schiffer–Tully scheme89 based on wavefunction overlap integrals between successive time steps. Electronic dynamics were integrated using 25 sub-steps per nuclear time step, and decoherence effects were included using the energy-based correction of Granucci et al.90 with a 0.1 Hartree parameter.
S2.4 Validation of the QM/MM scheme
The definition of the QM region is a critical factor for reliably capturing excited-state (ES) dynamics, particularly for proton transfer (PT) processes, which are known to be sensitive to the QM/MM boundary in flexible or solvated systems.96, 97 Although we focus on crystals, where structural rigidity simplifies this choice6, we assessed the robustness of our conclusions by expanding the QM region in the LLS system to include a third SO counterion. In our default setup, only one C-terminus of the Lys dimer interacts with a QM-level sulfate ion, while the second interfaces with a classical one. This modified setup allows us to evaluate whether explicitly treating both counterions at the QM level significantly alters the key vibrational modes involved in non-radiative decay. To this end, we selected five equilibrated initial conditions and re-equilibrated the system in the ground state for 1 ps using the enlarged QM region, following the same protocol as the second equilibration step described above. The final nuclear positions and velocities of these new QM/MM trajectories were then used to carry out 1 ps NAMD trajectories. All NAMD parameters were kept identical to those employed in the main simulations.
S3 Supplementary Text
S3.1 Determination of the Critical Aggregation Concentration
@@
Lys samples for the experimental morphological and optical characterization were prepared by dissolving the amino acid powder in double distilled water at the concentration of 1 g/mL (6.8 mol/L), which corresponds to the maximum of its solubility. Under these conditions, an increase of the volume of was observed, and the resulting solution exhibited a slight yellow coloration. The pH of the resulting solution was basic (pH = 11.8). Given the known ability of single amino acids to self-assemble into supramolecular structure98, the formation of Lys aggregates was investigated using two environment-sensitive fluorescent probes: ANS and pyr, as shown in Fig. S199. Both dyes are known to be able to change their fluorescence properties as function of the surrounding chemical environment (hydrophobic or hydrophilic). ANS shows a characteristic emission peak between 460-480 nm in hydrophobic environments, such as the inner micellar compartment100 or aromatic peptide aggregate spine101. The fluorescence of pyr, on the other hand, is characterized by five vibrational bands, where the intensity ratio between the first (I1 at 373 nm) and third (I3 at 383 nm) bands serves as a sensitive indicator of environmental polarity.
@@
The CAC of Lys was estimated from the inflection points of fluorescence titration curves (Fig. S1). CAC values determined using ANS (0.915 mol/L) and pyr (0.888 mol/L) fluorophores are in good agreement, supporting the reliability of the measurements. These values are slightly higher than the previously reported estimate of 0.5 mol/L by Stagi et al. 32, who identified a change in the fluorescence intensity slope as a function of concentration: at lower concentrations, intensity increased linearly, while at higher concentrations (from 0.4 mol/L), the rate of increase slowed, suggesting the formation of larger, less fluorescent aggregates. The fact that the ANS and pyr calculated values are below the Lys solubility (solubity/CAC 7.5) confirms that Lys aggregates, or smaller clusters, are present under the experimental conditions investigated.
S4 Supplementary Figures

### v2

@@ 4 Conclusions
The intrinsic fluorescence of lysine occurs only in the aggregated state while is absent at dilute concentrationsstagi2022. Therefore, it is essential to verify that experimental conditions are above the critical aggregation concentration (CAC). In this work, the CAC value of Lys in water was determined by using fluorescence spectroscopy with two established probe molecules: (a) 8-anilino-1-naphthalene sulfonic acid ammonium salt (ANS) and (b) pyrene (pyr). Fluorescence spectra were recorded at room temperature on a Jasco spectrofluorophotometer (Model FP-8550) using a quartz cuvette (1.0 cm path length). Experimental parameters were: excitation and emission bandwidths = 5 nm; recording speed = 120 nmmin-1, excitation wavelength = 350 nm (ANS) and 335 nm (pyr). For titrations, increasing aliquots of Lys solution were added to 200 L of 200 M ANS or 10 M pyr aqueous solutions. Blank spectra were subtracted, and fluorescence intensities were corrected for dilution. Titrations were performed in triplicate, and trends were analyzed by least-squares fitting. In the ANS assay, the CAC value was determined from a linear least-squares fit of the fluorescence intensity at 475 nm (excitation at 350 nm) versus Lys concentration, following Balasco et al.Balasco2024. In the pyr assay, the CAC was determined from the break-point of the fitting of the I1/I3 ration (excitation at 335 nm) versus Lys concentrationMANJU2011318.
S1.1.2 CAC Measuraments
@@
L-Lysine (Lys) samples for the experimental morphological and optical characterization were prepared by dissolving the amino acid powder in double distilled water at the concentration of 1 g/mL (6.8 mol/L), which corresponds to the maximum of its solubility. Under these conditions, an increase of the volume of was observed, and the resulting solution exhibited a slight yellow coloration. The pH of the resulting solution was basic (pH = 11.8). Given the known ability of single amino acids to self-assemble into supramolecular structureGOUR2021154, the formation of Lys aggregates was investigated using two environment-sensitive fluorescent probes: ANS and pyr, as shown in Fig. S2diaferia_2015. Both dyes are known to be able to change their fluorescence properties as function of the surrounding chemical environment (hydrophobic or hydrophilic). ANS shows a characteristic emission peak between 460-480 nm in hydrophobic environments, such as the inner micellar compartmentVaccaro_2007 or aromatic peptide aggregate spineDiaferia_2016. The fluorescence of pyr, on the other hand, is characterized by five vibrational bands, where the intensity ratio between the first (I1 at 373 nm) and third (I3 at 383 nm) bands serves as a sensitive indicator of environmental polarity.
@@
The CAC of Lys was estimated from the inflection points of fluorescence titration curves (Fig. S2). CAC values determined using ANS (0.915 mol/L) and pyr (0.888 mol/L) fluorophores are in good agreement, supporting the reliability of the measurements. These values are slightly higher than the previously reported estimate of 0.5 mol/L by Stagi et al. stagi2022, who identified a change in the slope of the lysine fluorescence intensity as a function of its concentration: at lower concentrations, intensity increased linearly, while at higher concentrations (from 0.4 mol/L), the rate of increase slowed, suggesting the formation of larger, less fluorescent aggregates. The fact that the ANS and pyr calculated values are below the Lys solubility (solubity/CAC 7.5) confirms that Lys aggregates, or smaller clusters, are present under the experimental conditions investigated.
The CAC was determined only under basic conditions (pH = 11.8), corresponding to the zwitterionic form of Lys, to confirm that the amino acid is aggregated under the experimental conditions used in this work. However, since pH strongly influences the structural and emissive properties of Lys, a more systematic investigation of the CAC as a function of pH would be an interesting direction for future studies.
S2 Supplementary Figures